SPECTROSCOPIC PROBES OF THERMAL EXCITATIONS IN LIQUID HELIUM

ERIK W. DRAEGER, Lawrence Livermore National Laboratory, P.O. Box 808, Livermore, California 94550 ; KEVIN K. LEHMANN, Department of Chemistry, Princeton University, Princeton, New Jersey 08544; \underline{ROGER E. MILLER}, Department of Chemistry, University of North Carolina at Chapel Hill, Chapel Hill, North Carolina 27599.

We have shown in previous studies that linear chains of HCN can be formed by cluster growth in helium nanodroplets. Although these clusters are linear, the parallel vibrational band s associated with the C-H stretches of many of them show strong Q branches, suggesting that the moment of inertia about the molecular axis is significant. The experimental spectra f or these systems are discussed in the context of recent theoretical calculations, using pat h integral methods, which show that the helium near the cluster can be thermally excited an d thus contribute to the moment of inertia about the axis of the cluster